CAT Chat Minutes
June 14, 2004
Information and Follow-up
Thomas Gog recalled on a past occasion that his team worked on their experiment for nearly one-half day before a floor coordinator told them that the control room had removed the shutter permit for a short time. He feels that they could have been told sooner, saving much undue work by the experimenters.
User answered by commending the APS for improving its communication and that he has been hearing the announcements as to the stability of the ring.
General Information
Machine Stats
Delivered beam: 400.00 hours
Percentage of scheduled time: 97.7%
Downtime during period: 9.40 hrs.
MTBF: 33.31 hrs.
Beamline Downtime: Summary by APS groups for period 5/25/2004 to 6/14/2004
DIN |
Date |
Beamline |
Status |
Duration |
Group |
Description |
Resolution |
395 |
05/25/04 |
13-BM |
Closed |
20 min |
AOD-IT |
GSE-CARS are having problems with certain PV's that allow them to receive PSS signals. Marty Smith is fixing the problem. |
ASD-CTL (Marty Smith) reset hardware to allow communications with the desired PVs . |
396 |
05/25/04 |
13-ID |
Closed |
58 min |
ASD-SI |
PS1 closes every time CAT tries to open PS2. |
ASD-SI Systems Engineer (J. Behrndt ) found device reporting "not ready". Shutter excercised several times and returned to proper operational status; limit switches properly indicating PS2 as being open. |
400 |
05/27/04 |
01-BM |
Closed |
46 min |
AOD-OAG |
Shutter permit removed "due to the BTS magnet being switched to a not to the dump state and the Top-Up key being disabled." per MCR shift log on Thursday May 27th, 2004, 15:00-23:00. MCR Downtime Record will record the first 30 minutes of downtime. (1-BM shutter closed for 9 minutes). |
Procedure Execution Manager (PEM) has been changed to remove PVs that should not be changed. Occurrence will not be repeated. |
403 |
05/26/04 |
22-BM |
Closed |
10 min |
ASD-SI |
M. Hahne reported PS1 and PS2 not showing open or closed. Device status showed position faults on both. Jon Hawkins is looking into the problem. |
FE-EPS system engineer (J. Behrndt ) discovered loose communications cable. Problem resolved when connector properly seated. |
405 |
05/28/04 |
13 |
Closed |
1 hr 20 min |
AOD-IT |
CAT firewall crashed. Reported by 13, 14, 23, 24 but probably affected most beamlines . |
System rebooted; network is functioning normally. MCR made an announcement regarding restoration of communications. |
406 |
06/04/04 |
18-ID |
Open |
1 hr 5 min |
Serious Fault #211 occurred (P8DMS2_NOSWITCH). Per FC standing order, FC reset the fault and called Roy Emerson. |
||
407 |
06/09/04 |
14-ID |
Closed |
1 hr 27 min |
ASD-SI |
BP reported problem with door lock on 14-ID-A Door #1 (manual). Minor fault #410 (A_DR1_CLSE_FAIL) occurred. PSS is investigating the problem. |
ASD-SI has realigned the door lock sensors so that the LEDs display the proper lock condition. |
408 |
06/09/04 |
18-ID |
Closed |
2 hr 23 min |
ASD-VAC |
Serious fault #69 (PS2_MIXUP-S) occurred. Per Roy Emerson, I reset the fault 3 times, and the fault re-occurred each time. I called Joe Gagliano , who came and discovered a leak at the solenoid valve on top of the storage ring. |
ASD-VAC will replace PS2 solenoid valve and rebuild the cylinder following scheduled access to Zone C on 06/15/2004. Refer to Work Request #15314. |
409 |
06/10/04 |
06-ID |
Closed |
30 min |
AOD-EOS |
O2 monitor alarmed reading 18.6%. FC measured readings with portable unit and found O2 level was normal at 20.9%. |
AOD-EOS (J. Ullian ) performed recalibration of O2 monitor to correct the long term drift of the sensor. |
410 |
06/10/04 |
18-ID |
Open |
20 min |
Mono shutter faulted shortly after opening the D shutter (located in station C) with PSS fault code 211 (P8DMS2_NOSWITCH). |
||
411 |
06/10/04 |
16-BM |
Closed |
1 hr 30 min |
ASD-SI |
16BM-B door close button is falling out of the PSS door box and is no longer functioning. |
ASD-SI System Engineer (R. Emerson) and technician (R. Gomez) responded to FC call in. Arrived at APS and effected repairs to door lock box hardware. Once repairs were completed the system was tested to insure proper functionality. |
412 |
06/11/04 |
19-BM |
Closed |
18 min |
ASD-SI |
Beamline recieved PSS fault #32 Chain B Dead. |
FC contacted ASD-SI on-call engineer and received permission to perform reset. Fault reset without incident. |
APS FY04 Long Range Operations Schedule (October 2003 - September 2004): Roger showed where we are on the operations schedule for top-up user operation in low emittance mode. Roger asked the group if there was a preference to have July 4th set aside for studies, beam run on the 5th. A user replied, why only one day that beam down, why not more? Roger answered that he did not know other than it was always done this way at the APS.